OpenStudy (anonymous):

Help Asap.... Medal and Fan given Which of the following equations best represents the graph of a function f(m)? f(y) = m f(x) = m y = f(m) x = f(m)

OpenStudy (anonymous):

@KlOwNlOvE @ikram002p

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@RaphaelFilgueiras

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

how did you get that ?

OpenStudy (anonymous):



OpenStudy (anonymous):

thank you do you mind helping me w/ a few other questions ?

OpenStudy (anonymous):

@RaphaelFilgueiras

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20